OpenStudy (anonymous):

What is the inverse of the equation y = 2x - 3?

OpenStudy (kainui):

The inverse of an equation is basically going backwards instead of from x to y, it's going from y to x. In order to reverse this, just switch the variables around and solve for your new "y" For example if your equation was y=2x then you would switch it to x=2y and solve for y which would be y=x/2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i think i have lost my mind let me try again suppose you had to solve \[2x-3=11\] what would you do? 1) ADD 3 2) DIVIDE by 2 inverse does the same thing \[f^{-1}(x)=\frac{x+3}{2}\]

OpenStudy (anonymous):

or you can write \[y=2x-3\] then switch \(x\) and \(y\) because that is what the inverse does, write \[x=2y-3\] then solve for \(y\) same two steps, same answer \[x=2y-3\] \[x+3=2y\] \[\frac{x+3}{2}=y\] \[f^{-1}(x)=\frac{x+3}{2}\]

OpenStudy (kainui):

Wait, are you correcting me because I'm wrong? I'm not sure... anymore...